OpenStudy (anonymous):

There are six students in the poetry club. One gets to be president of the club, another vice president, and another secretary. How many ways can these three positions be filled by the members of the club?

OpenStudy (kropot72):

The order of selection matters, therefore the number of permutations of the six students taken three at a time is required. Number of ways of filling three positions = 6P3 = 6!/3!

OpenStudy (kropot72):

\[\frac{6!}{3!}=\frac{6\times 5\times 4\times 3\times 2\times 1}{3\times 2\times 1}=you\ can\ calculate\]